OpenStudy (anonymous):

Jared Solved the equation below. What mistake did he make? 3b=25 3b-3=25-3 b=22

OpenStudy (anonymous):

it should be 3b/3

OpenStudy (anonymous):

and 25/3

OpenStudy (anonymous):

3b/3 = 25/3 b = 25/3